# arXiv:2602.23440v1[cs.CL] 26 Feb 2026

### v1

@@ A.3 Proof of Theorem 1
Part 1 shows the truncated estimator is never worse; we now show it can be $T$ times better. The intuition is that the total trajectory variance $\text{Var}[R(\tau)]$ is the sum of variances from all $T$ steps, but the truncated estimator only “sees” the variance from one step—hence the $1/T$ factor.
Under Assumption 2 (conditional independence), for each step $t$, the reward $r_{t}$ depends only on $a_{t}$ and $\tau_{<t}$, and future actions do not affect past rewards. Under this condition, the trajectory reward variance decomposes as:
$\text{Var}[R(\tau)]$ $=\text{Var}\!\left[\sum_{t=1}^{T}r_{t}\right]=\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]+\text{inter-step terms}.$
Under Assumption 2, the inter-step covariance terms are non-negative: conditional independence ensures that $\text{Cov}(r_{t},r_{t^{\prime}}\mid\tau_{<t})\geq 0$ for $t<t^{\prime}$, since any remaining dependence flows only through the sequential prefix structure. Therefore:
@@
A.4 Sample Efficiency
@@ A.5 Proof of Proposition 2
We now count total tokens generated. In standard GRPO, we sample $G$ complete trajectories of average length $L$, costing $G\cdot L$ tokens. In our method, each truncated sample generates only one step’s worth of tokens—approximately $L/T$ tokens, since the full trajectory’s $L$ tokens are spread over $T$ steps—and we repeat this sampling at each of the $T$ steps. The total cost is therefore:
@@
[…]
A.6 Variance Reduction and Convergence
@@
A.7 Bias-Variance Trade-off
@@
A.8 Early-Termination Bonus Details
@@
The bonus term $\lambda\cdot(B-t)/B$ in Eq. 6 encourages the model to produce an answer as soon as it has gathered sufficient information. Without such a term, the model may learn to issue superfluous search queries—each receiving a neutral or mildly positive reward from the LLM judge—even when the information needed to answer has already been retrieved. The bonus is largest when the model answers early (e.g., $\lambda\cdot\frac{3}{4}$ at step t=1t{=}1 with B=4B{=}4) and zero when it exhausts the full budget (t=Bt{=}B), creating a progressive incentive to terminate sooner. Crucially, this bonus only applies to answer candidates, so at any step where some of the $k$ sampled actions produce an answer and others produce a search query, the answer candidates receive a higher reward, creating a meaningful advantage signal that survives the group normalization in Eq. 4. This ensures the policy gradient directly reinforces early termination when additional search steps are unlikely to improve the answer.
A.9 Effect of Group Size $k$
@@
A.10 LLM-as-Judge Reward Prompts
We provide the exact prompts used for the three LLM-as-judge reward components described in Section 3.3. In each prompt, the placeholders in curly braces (e.g., {context}, {thinking}) are filled with the corresponding trajectory content at evaluation time. The judge is instructed to produce a chain-of-thought explanation before the score, enclosed in XML-style tags.

### v2

@@ A.3 Proof of Theorem 1
Part 1 shows the truncated estimator is never worse; we now show it can be $T$ times better. The intuition is that the total trajectory variance $\text{Var}[R(\tau)]$ is the sum of variances from all $T$ steps, but the truncated estimator only “sees” the variance from one step, hence the $1/T$ factor.
Under Assumption 2 (conditional independence), for each step $t$, the reward $r_{t}$ depends only on $a_{t}$ and $\tau_{<t}$, and given the prefix $\tau_{<t}$, the step reward $r_{t}$ is independent of rewards at other steps conditioned on their respective prefixes. We establish the desired inequality by applying the law of total variance recursively. For any step $t$, the law of total variance gives:
$\text{Var}[R(\tau)]=\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right]+\text{Var}_{\tau_{<t}}\!\left[\mathbb{E}[R(\tau)\mid\tau_{<t}]\right]\geq\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right].$ (18)
From Part 1 (using Assumption 1), we already have $\text{Var}[r_{t}\mid\tau_{<t}]\leq\text{Var}[R(\tau)\mid\tau_{<t}]$ for each prefix $\tau_{<t}$. Summing over all steps:
$\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]\leq\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right].$ (19)
Under Assumption 2, the per-step conditional variances account for distinct, non-overlapping sources of randomness (each step’s action is the sole source of variation given its prefix). Therefore:
@@
A.4 Proof Remarks
Remark 1 (Bias-Variance Trade-off in the Estimator Comparison).
Theorem 1 compares the variance of two different advantage estimators: the trajectory-level $\hat{A}_{i}$ (based on $R(\tau)$) and the step-level $\hat{A}_{t}^{(j)}$ (based on $r_{t}$). Because the truncated estimator targets the step-level reward rather than the full return, it is not an unbiased estimator of the trajectory-level advantage. The practical validity of this substitution rests on the step-level reward being a low-bias proxy for the contribution of action $a_{t}$ to the trajectory outcome, a condition favored by our setting’s short horizons ($T\leq 4$), externally grounded evaluation (retrieved documents are directly observable), and evaluative (rather than predictive) reward design (Section 3.3). We discuss this trade-off further in Appendix A.8.
Remark 2 (Credit Assignment).
A.5 Sample Efficiency
@@ A.6 Proof of Proposition 2
We now count total tokens generated. In standard GRPO, we sample $G$ complete trajectories of average length $L$, costing $G\cdot L$ tokens. In our method, each truncated sample generates only one step’s worth of tokens, approximately $L/T$ tokens, since the full trajectory’s $L$ tokens are spread over $T$ steps, and we repeat this sampling at each of the $T$ steps. The total cost is therefore:
@@
The two sources of savings, $T\times$ fewer samples needed (due to lower per-sample variance) and $T\times$ fewer tokens per sample (due to truncation), together yield a $T^{2}$ reduction; paying back one factor of $T$ for repeating across all $T$ steps gives a net $T$-fold improvement over the $G\cdot L$ tokens required by standard GRPO. ∎
A.7 Variance Reduction and Convergence
@@
A.8 Bias-Variance Trade-off
@@
A.9 Limitations
Our truncated sampling constructs a single trajectory by greedily extending one selected action per step, which limits exploration compared to maintaining multiple full trajectories. While the short horizons ($T\leq 4$) and externally grounded rewards in our setting mitigate this, the approach may be less effective in domains with longer horizons or where locally promising actions frequently lead to dead ends. Additionally, our method relies on a large LLM judge (Gemma3-27B) for dense supervision, which introduces computational overhead and makes the overall improvement dependent on the quality of the judge model.
A.10 Early-Termination Bonus Details
@@
The bonus term $\lambda\cdot(B-t)/B$ in Eq. 6 encourages the model to produce an answer as soon as it has gathered sufficient information. Without such a term, the model may learn to issue superfluous search queries, each receiving a neutral or mildly positive reward from the LLM judge, even when the information needed to answer has already been retrieved. The bonus is largest when the model answers early (e.g., $\lambda\cdot\frac{3}{4}$ at step t=1t{=}1 with B=4B{=}4) and zero when it exhausts the full budget (t=Bt{=}B), creating a progressive incentive to terminate sooner. Crucially, this bonus only applies to answer candidates, so at any step where some of the $k$ sampled actions produce an answer and others produce a search query, the answer candidates receive a higher reward, creating a meaningful advantage signal that survives the group normalization in Eq. 4. This ensures the policy gradient directly reinforces early termination when additional search steps are unlikely to improve the answer.
A.11 Why Process Rewards Succeed in Search but Not Math
While process reward models have proven effective as verifiers for mathematical reasoning (Lightman et al., 2023), using step-level rewards as RL training signals has generally underperformed outcome-based verifiable rewards for policy optimization in math (Uesato et al., 2022). We argue that search-augmented reasoning possesses structural properties that make process rewards fundamentally more reliable as RL training signals, and that our truncated sampling design addresses the remaining risks.
External Grounding Eliminates Capability Mismatch.
In math, a process reward model (PRM) may assign high reward to an elegant proof strategy that the actor model cannot complete, for example, a sophisticated algebraic manipulation that only a larger model could carry out. This “capability mismatch” means the PRM rewards steps that look locally promising but lead to dead ends. In search, the “hard part” of each step is outsourced to the search engine: the model only needs to formulate a well-targeted query, and the retrieval system handles the actual information lookup. The capability ceiling for writing a good query is much lower than for completing an advanced proof, so the process reward cannot favor steps that exceed the actor’s abilities.
Evaluative vs. Predictive Rewards.
[…]
Short Horizons and Natural Decomposability.
Search trajectories in our setting have $T\leq 4$ steps, each consisting of a discrete think-then-query action with a natural boundary (the search engine call). Math proofs can involve dozens of tightly coupled steps where the value of step 5 depends critically on steps 15–20. With shorter, more modular trajectories, the gap between local step quality and trajectory-level outcome is inherently smaller, reducing the risk that process rewards mislead the policy.
Truncated Sampling Prevents Local Minima.
Even granting that search rewards are more locally informative, standard full-trajectory sampling still suffers from prefix confounding: a locally good step on a bad trajectory gets penalized, and a mediocre step on a lucky trajectory gets rewarded. This is exactly how local minima arise in PRM-based RL, the policy gets “trapped” by rewarding steps that merely co-occur with good outcomes. Our truncated sampling eliminates this confound: all $k$ candidates share the same prefix $\tau_{<t}$, so advantages reflect only the current action’s quality. The reward-weighted sampling strategy (temperature $\eta=0.7$) further prevents greedy collapse into local optima during trajectory construction. The ablation in Table 2 provides direct evidence: variant (a), which uses LLM-judge rewards without truncated sampling, underperforms full Slate most on the hardest benchmarks, precisely where prefix-level confounds are most harmful.
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
Direct 0.152 0.434 0.178 0.145 0.249 0.018 0.030 0.172
CoT 0.168 0.451 0.196 0.163 0.257 0.032 0.072 0.191
RAG 0.312 0.504 0.296 0.251 0.221 0.051 0.076 0.244
IRCoT 0.286 0.478 0.268 0.171 0.196 0.059 0.245 0.243
Search-o1 0.267 0.465 0.254 0.240 0.207 0.045 0.316 0.256
SFT 0.301 0.491 0.286 0.191 0.248 0.039 0.112 0.238
R1 (no search) 0.258 0.467 0.243 0.191 0.277 0.051 0.232 0.246
ZeroSearch 0.348 0.531 0.365 0.260 0.234 0.056 0.096 0.270
ReSearch 0.339 0.518 0.342 0.261 0.228 0.074 0.184 0.278
Search-R1 0.406 0.587 0.435 0.284 0.273 0.049 0.088 0.303
StepSearch 0.329 0.339 0.181 0.328
Slate (Ours) 0.425 0.603 0.451 0.351 0.368 0.216 0.361 0.396
A.12 Effect of Group Size $k$
@@
A.13 LLM-as-Judge Reward Prompts
We provide the exact prompts used for the three LLM-as-judge reward components described in Section 3.3. In each prompt, the placeholders in curly braces are filled with the corresponding trajectory content at evaluation time. The judge is instructed to produce a chain-of-thought explanation before the score, enclosed in XML-style tags.